OpenStudy (anonymous):

When you have 13pi/3 -2pi , why do you get 7pi/3? Wouldn't it be 11pi/3? I am so confused.

OpenStudy (amistre64):

what is 13/3 - 2 ?

OpenStudy (amistre64):

since 3/3 = 1, this can be written as: 13/3 - 3/3 - 3/3 (13 - 3- 3)/3

OpenStudy (anonymous):

i see where you got the 7 from I just understand why you did 3/3. Does it have something to do with having to have the same denominators in order to subtract?

OpenStudy (amistre64):

its simpler for us to work the subtraction yes. common denominators place all the values we are to work with on a common playing field. 13/3 and 2/1 , i say are not wearing the same pants and we have to redress them

OpenStudy (amistre64):

13/3 is not equal to 13 13-2 = 11, is what you prolly saw in your mind; but by doing that you make up your own problem by changing 13/3 into 13

OpenStudy (anonymous):

So what happend to the 2pi in the midst of all this?

OpenStudy (amistre64):

pi is pi is pi ... we factored it out and worked the fractions

OpenStudy (amistre64):

\[\frac abx+\frac mnx=x(\frac ab+\frac mn)\]

OpenStudy (anonymous):

so then it becomes 1 over 1 which you multiply by 3 to get the 3/3 to have the same denominators so you can subtract?

OpenStudy (amistre64):

\[\frac{13}{3}-2\]\[1=\frac 33\] \[-1=-\frac 33\] \[-2=-1-1=-\frac 33 - \frac 33\] \[\frac{13}{3}-2=\frac {13}3-\frac 33-\frac 33\]

OpenStudy (amistre64):

OR, yeah, we multiply 2 by 3/3

OpenStudy (amistre64):

\[\frac{13}{3}-2(1)=\frac{13}{3}-2(\frac 33)\] \[\frac {13}3-\frac 63\]

OpenStudy (anonymous):

ok thank you so much! i think i understand it now! i really appreciate your help.

OpenStudy (amistre64):

good luck